OpenStudy (luigi0210):

Differentiate the following:

OpenStudy (luigi0210):

\[\LARGE f(x)=(x+x^{-1})^3\]

OpenStudy (ikram002p):

\(\Huge \color{green}{(\frac{x^2+1}{x})^3}=(x+x^-1)^3\)

OpenStudy (ikram002p):

\(\Huge \text{let} \frac{x^2+1}{x} =u\)

OpenStudy (ikram002p):

then u wud have \(\Huge (u^3)'=3u'(u)^2\)

OpenStudy (anonymous):

Why not do \[ \sqrt[3]{f(x)} = x+x^{-1} \]And then differentiate?

OpenStudy (ikram002p):

its to power 3 not 2

OpenStudy (anonymous):

You would get:\[ [f(x)]^{-2/3}f'(x) = 1+\ln(x) \]

OpenStudy (anonymous):

So \[ f'(x) = \left(1+\ln(x)\right)\left(\left(x+x^{-1}\right)^{3}\right)^{2/3} =\left(1+\ln(x)\right)\left(x+x^{-1}\right)^{2} \]

OpenStudy (ikram002p):

i wud get 3(1-x^-2)(x+x^-1)^2

OpenStudy (ikram002p):

im a bit confused nw :o

OpenStudy (anonymous):

Hmmm, maybe I differentiated wrong.

OpenStudy (ikram002p):

i dnt know were did u came with the ln formula its derevative not integral even if u let it f(x)=(x+x^-1)^3 let u =x+x^-1 then u'=1-x^-2 and (u^3)'=3u'u^2=3(1-x^-2)(x+x^-1)^2

OpenStudy (anonymous):

probably should be \(x^{-2}\) instead of \(\ln(x)\)

OpenStudy (ikram002p):

(x^-1)'=-x^-2 (ln x)'=1/x :o

OpenStudy (anonymous):

This is so easy, but I can't help my student Luigi.